OpenStudy (anonymous):

Prove the identity. (tanx+1)^2=(1+2sinxcosx)/(cos^2x) please help! so far, i have (sin^2x/cos^2x)+(2sinx+1/cosx) I'm not sure what do next! it seems wrong!

OpenStudy (asnaseer):

You are virtually there - just add the terms you got to

OpenStudy (asnaseer):

You have two fractions, one with denominator \(\cos^2(x)\) and one with denominator of \(\cos(x)\). So the common denominator is \(\cos^2(x)\).

OpenStudy (cwrw238):

(tanx+1)^2 = tan^2x + 2tanx + 1 sin^2x / cos^2x + 2sinx/cosx + 1 = [sin^2x + 2sinxcosx + cos^2x] / cos^2x sin^x + cos^x = 1 so :

OpenStudy (anonymous):

OH i see what i did wrong. THANK YOU SO MUCH :D

OpenStudy (cwrw238):

your welcome